# FENDA-FL: Personalized Federated Learning on Heterogeneous Clinical Datasets

### v2

@@ 3.3 Federated Checkpointing and Evaluation
flamby_results_table Fed-Heart-Disease: Acc. Fed-IXI: Dice Fed-ISIC2019: Balanced Acc. Global Local Global Local Global Local FedAvg $0.724$ ($0.000$) $0.724$ ($0.000$) $0.9845^{*}$ ($0.0001$) $0.9846^{*}$ ($0.0001$) $\underline{0.657}$ ($0.004$) $0.644$ ($0.004$) FedAdam $0.719$ ($0.000$) $0.742$ ($0.000$) $0.9811^{*}$ ($0.0002$) $0.9811^{*}$ ($0.0002$) $0.647$ ($0.013$) $0.637$ ($0.012$) FedProx $0.716$ ($0.000$) $0.721$ ($0.000$) $0.9842^{*}$ ($0.0001$) $0.9844^{*}$ ($0.0001$) $0.633$ ($0.022$) $0.652$ ($0.022$) SCAFFOLD $0.711$ ($0.000$) $0.682$ ($0.000$) $0.9827^{*}$ ($0.0000$) $0.9827^{*}$ ($0.0000$) $\mathbf{0.672}^{*}$ ($0.013$) $0.669$ ($0.010$) MOON $0.729$ ($0.009$) $0.741$ ($0.009$) $0.9852^{*}$ ($0.0001$) $0.9851^{*}$ ($0.0002$) $0.620$ ($0.019$) $0.604$ ($0.017$) FedPer $\underline{0.814}^{*}$ ($0.002$) $0.9852^{*}$ ($0.0001$) $0.634$ ($0.014$) APFL $0.801^{*}$ ($0.006$) $\mathbf{0.9864}^{*}$ ($0.0002$) $0.608$ ($0.011$) PerFCL $0.805^{*}$ ($0.018$) $0.9847^{*}$ ($0.0001$) $0.584$ ($0.019$) FENDA-FL $\mathbf{0.815}^{*}$ ($0.000$) $\underline{0.9856}^{*}$ ($0.0001$) $0.607$ ($0.011$) Silo $0.748$ ($0.000$) $0.9737$ ($0.0040$) $0.634$ ($0.025$) Central $0.732$ ($0.000$) $0.9832$ ($0.0002$) $0.683$ ($0.020$) Client 0 $0.735$ ($0.000$) $0.9648$ ($0.0123$) $0.549$ ($0.014$) Client 1 $0.649$ ($0.000$) $0.9722$ ($0.0004$) $0.387$ ($0.029$) Client 2 $0.589$ ($0.000$) $0.9550$ ($0.0015$) $0.554$ ($0.010$) Client 3 $0.629$ ($0.000$) $0.444$ ($0.016$) Client 4 $0.326$ ($0.018$) Client 5 $0.285$ ($0.039$)
Table 2: Performance metrics for the FLamby tasks. Values in parentheses are $95$% confidence-interval (CI) radii. Global and local refer to the checkpointing strategies and Silo metrics are measured as described in Section 3.3. Bold denotes the best performing federated strategy, while underlined values are second best. Asterisks imply a value is better than siloed training considering CIs. Note that each task has a different number of clients. For example, Fed-Heart-Disease has only four.
4 Results
In this section, all experiments use the checkpointing and evaluation strategies detailed in Section 3.3. Trainable parameter counts for the models are documented in Appendix C. An effort is made to maintain similar numbers of parameters across approaches. For the FLamby experiments, $15$ server rounds are performed for Fed-Heart-Disease and Fed-ISIC2019, while $10$ are performed for Fed-IXI. Each client completes $100$ local training steps within each round. Training settings for centralized and local training for each dataset are discussed in Appendix A. The loss functions, batch sizes, models, and metrics are all chosen to match the FLamby benchmark settings, unless otherwise specified. For both GEMINI tasks, $50$ server rounds are performed. The clients in Mortality and Delirium prediction complete $2$ and $4$ local epochs, respectively, at each server round. All APFL training initializes $\alpha=0.5$. GEMINI training setups are detailed in Appendix B.
4.1 FLamby: Fed-Heart-Disease
The baseline model for Fed-Heart-Disease is logistic regression. The feature space is only $13$ dimensions. Hence, this model incorporates only $14$ parameters. However, a deeper model is required for several approaches to implement the strategy. Both MOON and FedPer use a single linear layer and ReLU activation for their feature extractors. For PerFCL and FENDA-FL, both feature extractors are also single linear layers with a ReLU activation. For each approach, the classification head is a linear layer followed by a sigmoid. For APFL, a two layer dense neural net (DNN) is used for both the global and local model with a ReLU in between followed by a sigmoid. Using two-layer DNNs with an equivalent number of parameters for the other strategies generally did not improve performance, as seen in Appendix D.
The results, displayed in Table LABEL:flamby_results_table, show that FENDA-FL is the best performing approach, followed closely by FedPer. Notably, all personalized FL methods outperform centralized training by a large margin. Furthermore, the personalized methods are the only FL mechanisms to beat siloed training, where each client trains and evaluates on its own data. The results also suggest that local checkpointing has some advantages for this task, except for SCAFFOLD. In Appendix F, ablation studies are carried out demonstrating the significant advantage of using federated checkpointing, especially local checkpointing, over a standard fixed server-round strategy.
4.2 FLamby: Fed-IXI
As a 3D-segmentation task, the metric for Fed-IXI is a Dice score. The model architecture is a 3D U-Net (Cicek et al., 2016) and the implementation in FLamby is used. The encoder produces voxel-wise features of the 3D volume at different resolutions that are concatenated along the channel dimension with features from the decoder of the same resolution. The per-voxel channel contents of features at the highest resolution are then used to produce a binary classification for each voxel. For a fair comparison of other approaches to APFL, PerFCL, and FENDA-FL, which require architecture modifications, the number of such channels in the U-Net is increased to $12$ for other methods, compared with the FLamby default of $8$. The models used for APFL, PerFCL, and FENDA-FL each produce $8$ channels. The PerFCL and FENDA-FL architectures use the U-Net without a classifier for both the global and local modules. The resulting activations are concatenated and fed into a classification layer. The FedPer and MOON models are split between the U-Net base and the voxel-wise binary classifier.
As seen in Table LABEL:flamby_results_table, each strategy produces a model that performs this task well. However, there is a clear advantage to central training and an even larger one from FL. Among non-personalized FL methods, models trained with FedAvg, FedProx, and MOON surpass central training. All personalized FL approaches also surpass the centrally trained model and produce four of the top five best metrics. APFL and FENDA-FL yield the best models, with APFL beating FENDA-FL by a small margin. Importantly, additional experiments, discussed in Appendix E, show that this gap can be closed with thoughtful architecture design for FENDA-FL. Finally, as with Fed-Heart-Disease, a small advantage is seen in using local checkpointing in most methods.
These results stand in contrast to those reported in the original FLamby benchmark. There, all FL strategies severely under-performed even local training. In that work, a standard SGD optimizer is used for client-side optimization, whereas we use AdamW. It should also be noted that the models for this task, along with those of Fed-ISIC2019, employ Batch Normalization layers. When using FedAdam, it is critical that stats tracking be turned off for these layers to avoid issues during the aggregation process. This phenomenon is discussed in Appendix H. We speculate that a related issue may have impeded the convergence of the FedOpt family in their experiments.
4.3 FLamby: Fed-ISIC2019
The model used for Fed-ISIC2019 is EfficientNet-B0 (Tan and Le, 2019), which has been pretrained on ImageNet, with a linear final layer. The classifier layer for MOON and FedPer is defined as the last linear layer and the remainder of the model is the feature extractor. For APFL, the full model is used for the global and local modules. Finally, the PerFCL and FENDA-FL models remove the final layer and use the remaining architecture as the local and global feature extractors. The classification head is a two-layer DNN with a ReLU between the layers. For a fair comparison, the first $13$ layers of EfficientNet-B0 are frozen for APFL, PerFCL, and FENDA-FL.
@@ 4.3 FLamby: Fed-ISIC2019
fed_isic_generalization Refer to caption
Figure 2: Generalization measurements for a selection of FL and non-FL approaches. Rows correspond to a model trained using either client-only data, centralized data, or with the indicated FL method. Columns denote the performance of that model on the local test data of the named client.
[…]
mortality_results_table Mean Accuracy Mean AUC/ROC Global Local Global Local FedAvg $0.8976$ ($0.0002$) $0.8971$ ($0.0006$) $0.8180^{*}$ ($0.0005$) $0.8177^{*}$ ($0.0005$) FedAdam $0.8954$ ($0.0000$) $0.8954$ ($0.0000$) $0.8128$ ($0.0005$) $0.8123$ ($0.0008$) FedProx $0.8957$ ($0.0004$) $0.8956$ ($0.0003$) $0.8174^{*}$ ($0.0004$) $0.8172^{*}$ ($0.0006$) SCAFFOLD $0.8955$ ($0.0003$) $0.8954$ ($0.0000$) $\mathbf{0.8186}^{*}$ ($0.0002$) $0.8184^{*}$ ($0.0002$) MOON $0.8968$ ($0.0008$) $0.8970$ ($0.0002$) $\underline{0.8185}^{*}$ ($0.0002$) $0.8181^{*}$ ($0.0008$) FedPer $\mathbf{0.8985}$ ($0.0005$) $0.8179^{*}$ ($0.0006$) APFL $0.8978$ ($0.0004$) $0.8138$ ($0.0017$) PerFCL $0.8979$ ($0.0008$) $0.8117$ ($0.0012$) FENDA-FL $\underline{0.8980}$ ($0.0004$) $0.8181^{*}$ ($0.0001$) Silo $0.8978$ ($0.0009$) $0.8140$ ($0.0012$) Central $0.8981$ ($0.0005$) $0.8193$ ($0.0005$) Client 0 $0.8966$ ($0.0014$) $0.8052$ ($0.0030$) Client 1 $0.8930$ ($0.0025$) $0.7657$ ($0.0064$) Client 2 $0.8943$ ($0.0060$) $0.7591$ ($0.0116$) Client 3 $0.8954$ ($0.0000$) $0.7994$ ($0.0034$) Client 4 $0.8955$ $(0.0005)$ $0.7974$ ($0.0037$) Client 5 $0.8973$ $(0.0014)$ $0.7878$ ($0.0043$) Client 6 $0.8954$ $(0.0000)$ $0.8019$ ($0.0035$)
@@
The results in Table LABEL:flamby_results_table show that all FL strategies perform better than local models, with SCAFFOLD providing the best results. While it does not surpass centralized training, it is quite close. Further, each of the non-personalized FL methods equals or surpasses siloed evaluation, though only SCAFFOLD’s model is statistically significant. As with Fed-IXI, these results improve upon those of the original FLamby benchmark, where none of the FL methods showed strong performance. This led the authors to conclude that federated training was not a useful approach for this task. The results presented here compellingly support the opposite conclusion.
For this task, each of the personalized FL approaches under-perform. While the results are better than any of the local models, non-personalized methods, with the exception of MOON, materially out-perform these techniques. Moreover, they fail to eclipse siloed accuracy. Some insight into where these approaches fall short is found in Figure LABEL:fed_isic_generalization. SCAFFOLD outperforms the personalized models on test data drawn from Client 1 and Client 2. This is reversed when considering Client 4. However, the most significant difference is that all personalized models perform quite poorly on the test data of Client 5. Of these models, FedPer provides the best results, but is still substantially worse than SCAFFOLD.
Notably, Client 5 has the smallest dataset by a large margin; see Table LABEL:fl_dataset_distributions in the appendix. Moreover, the data appears to be of poor quality, with locally trained models only achieving a balanced accuracy of $0.425$ on the test set. Alternatively, models trained only on data from Client 0 generalize well to Client 5. We hypothesize that the low quality of Client 5’s data impedes personalized FL model training. In future work, we aim to consider ways of overcoming this issue. Some potential avenues include additional loss terms to ensure the global and local feature extractors remain individually useful for classification or using a global FL approach to initialize model components as a warm start for further fine-tuning. Interestingly, for this task, local checkpointing appears disadvantageous compared to global checkpointing.
[…]
delirium_results_table Mean Accuracy Mean AUC/ROC Global Local Global Local FedAvg $0.7987$ ($0.0127$) $0.8025$ ($0.0130$) $0.8302^{*}$ ($0.0146$) $0.8274^{*}$ ($0.0078$) FedAdam $0.7689$ ($0.0036$) $0.7688$ ($0.0127$) $0.7897$ ($0.0078$) $0.7881$ ($0.0055$) FedProx $\mathbf{0.8095}^{*}$ ($0.0036$) $0.8056$ ($0.0037$) $0.8488^{*}$ ($0.0043$) $\underline{0.8504}^{*}$ ($0.0042$) SCAFFOLD $0.7480$ ($0.0000$) $0.7480$ ($0.0000$) $0.5491$ ($0.0469$) $0.5491$ ($0.0469$) MOON $0.8035$ ($0.0103$) $\underline{0.8085}$ ($0.0086$) $0.8310^{*}$ ($0.0100$) $0.8263^{*}$ ($0.0077$) FedPer $0.8082^{*}$ ($0.0024$) $0.8462^{*}$ ($0.0035$) APFL $0.8031$ ($0.0052$) $0.8430^{*}$ ($0.0062$) PerFCL $0.7893$ ($0.0077$) $0.8111$ ($0.0089$) FENDA-FL $0.8064$ ($0.0036$) $\mathbf{0.8518}^{*}$ ($0.0017$) Silo $0.7936$ ($0.0102$) $0.8037$ ($0.0072$) Central $0.8114$ ($0.0068$) $0.8458$ ($0.0026$) Client 0 $0.7653$ ($0.0070$) $0.7977$ ($0.0128$) Client 1 $0.7781$ ($0.0063$) $0.7795$ ($0.0161$) Client 2 $0.7805$ ($0.0036$) $0.8185^{*}$ ($0.0021$) Client 3 $0.7634$ ($0.0062$) $0.7539$ ($0.0055$) Client 4 $0.7924$ ($0.0061$) $0.8179$ ($0.0076$) Client 5 $0.8028$ ($0.0040$) $0.8429^{*}$ ($0.0018$)
@@
4.4 GEMINI: Mortality
For mortality prediction, the goal is to predict in-hospital death, defined as mortality status at hospital discharge. A collection of electronic-health-record and administrative features, such as interventions, imaging descriptions, and blood transfusion data, are used to train a DNN. Except for APFL, PerFCL and FENDA-FL, all approaches, including locally and centrally trained models, use the same DNN architecture with five layers. MOON and FedPer use the first four layers for feature extraction and the final layer as a classification head. APFL, PerFCL, and FENDA-FL require different structures, but the number of parameters remains equal. In APFL, the same architecture is used but with half of the parameters for each of the global and local components. For FENDA-FL and PerFCL, the global and local modules each have two layers with half the parameters of the full-size model. The classifier module has three layers.
The performance of mortality models is assessed based accuracy and AUC/ROC. The results are shown in Table LABEL:mortality_results_table. FL methods improve upon locally trained models, significantly reducing the performance gap to the centrally trained model. This is especially true when considering local model performance measured via AUC/ROC.
@@ 4.4 GEMINI: Mortality
In terms of accuracy, FedAvg is the only standard FL method that rivals the personalized FL algorithms and approaches siloed evaluation. On the other hand, most methods surpass the siloed AUC/ROC evaluation metric, nearing centrally trained model performance. Specifically, FedProx, SCAFFOLD, and MOON beat siloed training based on AUC/ROC, but did not in terms of accuracy. This may be due to the class imbalance in this dataset. Considering both measures, FedPer and FENDA-FL provide the best results, both surpassing the siloed evaluation metrics. Each method is competitive with non-personalized approaches in terms of AUC/ROC and outperform them by a large margin in terms of accuracy.
4.5 GEMINI: Delirium
The goal of this task is to determine whether delirium occurred during a hospital stay in order to facilitate quality measurement and benchmarking, because delirium is a leading cause of preventable harm in hospitals. Labeling delirium cases is labor-intensive, requiring specialized medical knowledge. Even in the GEMINI consortium, only a relatively small population has been labeled, see Table LABEL:fl_dataset_distributions. Features such as lab measurements, medications, and diagnosis are used to perform binary classification. More details are available in Appendix B.2.
A DNN with six layers is used for all non-personalized FL methods. MOON and FedPer use the first five layers as a feature extractor and the final layer as the classifier. APFL’s global and local components are each half the size of the full-sized model. PerFCL uses two five-layer DNNs as the local and global feature extractors and a final linear layer as the classifier. FENDA-FL’s model consists of a two-layer and a four-layer feature extractor as the local and global modules, respectively. The classifier is a linear transform. The number of trainable parameters is kept approximately constant across all methods, as documented in Appendix C. Note that the FENDA-FL network is asymmetric. Appendix E discusses experimental results associated with various architecture choices for this task and Fed-IXI.

### v3

@@ 3.4 Federated Checkpointing and Evaluation
Figure 1: Feature distributions for patient gender (a), age (b), and length of stay at care units (c) for the Mortality dataset and units per medication group (d), sedative units (e), and number of room transfers per patient (f) in the Delirium dataset. Note that some features are normalized and smoothed for visualization.
4 Clinical Datasets
In this work, five distinct clinical datasets are considered. Three are drawn from the FLamby benchmark (du Terrail et al., 2022a) and two are privately held within the GEMINI consortium (Verma et al., 2017). Each represents a clinically relevant task and data derived from real patients. The selected FLamby datasets are Fed-Heart-Disease which is a binary task predicting whether a patient has heart disease from a collection of measurements, Fed-IXI which considers MR images of patients’ heads with the target of 3D segmentation of brain-tissue, and Fed-ISIC2019 which is a multi-class dermatology dataset consisting of 2D images. The task is to classify the images into one of eight different melanoma categories. More details on these datasets are found in Appendix D.
The GEMINI datasets consist of clinical data from seven hospitals in Canada. The data is anonymized by removing patient identifiers, yet still contains highly sensitive patient-level information. As such, some dataset details are provided in the sections below and Appendix E, while other information is withheld in accordance with the GEMINI data sharing and research ethics protocols. The GEMINI tasks are critical to improving patient care and are distinct from the FLamby tasks in several ways. Predictions are based on tabular data with a large number of features. Clients have access to a sizeable number of labeled data points. Finally, label imbalance in both settings increases training complexity. Due to the sensitive nature of the data and GEMINI’s closed environment, the datasets and code are not publicly released. However, they are available to researchers with access to GEMINI. [^2] GEMINI access may be requested at https://geminimedicine.ca/access-gemini-data/
4.1 GEMINI: Mortality
This task is in-hospital mortality prediction, which determines the probability of a patient’s latest encounter ending in mortality. The cohort consists of patients admitted to general internal medicine (GIM) wards and discharged between 2015 to 2020, prior to COVID-19, at seven large hospitals. De-identified electronic health records (EHR) and administrative information from a total of $114,662$ patients are used for training, with $28,665$ data points reserved for testing across all hospitals. Data heterogeneity across hospitals for a selection of features is visualized in Figures 1(a)–1(c). Additional details and patient characteristics for a selection of features are described in Appendix E.1. Note that related heterogeneity visualizations for the FLamby datasets appear in the original paper.
4.2 GEMINI: Delirium
The goal of this task is to determine whether delirium occurred during a hospital stay in order to facilitate quality measurement and benchmarking, because delirium is a leading cause of preventable harm in hospitals. Labeling delirium cases is labor-intensive, requiring specialized medical knowledge. Even in the GEMINI consortium, only a relatively small population has been labeled, see Table 8. The total cohort consists of $5,770$ admissions including EHR and administrative features of patients admitted to GIM wards from 2010 to 2019 across six hospitals. Features such as lab measurements, medications, and diagnosis are used to perform binary classification. Figures 1(d)–1(f), exhibit heterogeneity for a subset of features. More details are available in Appendix E.2.
5 Experiments and Results
In this section, results associated with three experiments are discussed. The main benchmarking and evaluation results for all datasets are detailed in Section 5.1. In Section 5.2, ablation studies are performed to provide a deeper analysis of the proposed checkpointing approaches. The third experiment, discussed in Section 5.3, considers the robustness of the FL approaches to extreme feature heterogeneity in the context of the GEMINI Delirium task. In addition, Appendix F reports a final architecture study of the FENDA-FL method highlighting its structural advantage over PerFCL and other approaches.
Throughout the discussion to follow, it is important to note that certain FL approaches require specific model modifications and choices with respect to their setup. For those interested, these changes are thoroughly detailed in Appendix B for each approach and dataset. As shown in Table 6, care has been taken to keep the number of trainable parameters approximately constant across different methods.
5.1 Main Results
The main results of this work are exhibited in this section. All experiments use the checkpointing and evaluation strategies detailed in Section 3.4. Table 1 provides a succinct representation of the overall results. Each method is ranked based on the value achieved for each task and metric considered in the experiments. For approaches with both global and local checkpoints, the metric value of the best performing checkpoint strategy is used to establish rank. Generally, it is evident that personalized approaches provide measurable benefits for the selected tasks over all baselines and non-personalized approaches. FENDA-FL often yields high-performing models and has the highest average rank across all measured metrics, followed closely by Ditto and FedPer. It is important to note that Ditto requires significantly more computational resources, compared with other methods, due to the need for two full-model backpropagation passes at each iteration. Client-side optimizers and tuned hyper-parameters for these experiments are outlined in Appendices C-E. Detailed results for each task are analyzed in the sections below.
Table 1: FL method rankings across all tasks and metrics considered. Non-personalized and personalized FL methods are separated by the horizontal line.
Fed-Heart Fed-IXI Fed-ISIC Mortality Delirium Avg. Rank
Metric Acc. Dice Bal. Acc. Acc. AUC/ROC Acc. AUC/ROC
FedAvg $8$ $7$ $3$ $6$ $5$ $6$ $6$ $5.86$
FedAdam $7$ $10$ $5$ $10$ $9$ $9$ $9$ $8.43$
FedProx $9$ $8$ $4$ $8$ $7$ $1$ $2$ $5.57$
SCAFFOLD $10$ $9$ $2$ $9$ $2$ $10$ $10$ $7.43$
MOON $6$ $4$ $7$ $7$ 3 $4$ $5$ $5.14$
FedPer $2$ $4$ $6$ $2$ $6$ $2$ $3$ $3.57$
Ditto $4$ $3$ $1$ $1$ $1$ $7$ $7$ $3.43$
APFL $5$ $1$ $8$ $5$ $8$ $5$ $4$ $5.14$
PerFCL $3$ $6$ $10$ $4$ $10$ $8$ $8$ $7.00$
FENDA-FL $1$ $2$ $9$ $3$ $4$ $3$ $1$ $\mathbf{3.29}$
5.1.1 FLamby results
For the FLamby experiments, $15$ server rounds are performed for Fed-Heart-Disease and Fed-ISIC2019, while $10$ are performed for Fed-IXI. Each client completes $100$ local training steps within each round. Training settings for centralized and local training for each dataset are discussed in Appendix D. The loss functions, batch sizes, models, and metrics are all chosen to match the FLamby benchmark settings, unless otherwise specified. The full set of results are displayed in Table 2.
For Fed-Heart-Disease, FENDA-FL is the best performing approach, followed closely by FedPer. In fact, all personalized FL methods outperform centralized training by a large margin. Furthermore, the personalized methods are the only FL mechanisms to significantly beat siloed training, where each client trains and evaluates on its own data. The non-personalized FL methods largely under-perform. Only the locally checkpointed MOON model produces a mean accuracy above the siloed metric. Finally, the results also suggest that local checkpointing has some advantages for this task, except for SCAFFOLD.
Table 2: Performance metrics for the FLamby tasks. Values in parentheses are $95$% confidence-interval (CI) radii. Global and Local refer to the checkpointing strategies and Silo metrics are measured as described in Section 3.4. Bold denotes the best performing FL strategy, while underlined values are second best. Asterisks imply a value is better than siloed training considering CIs. Note that the number of clients varies for each task. For example, Fed-Heart-Disease has only four.
Fed-Heart-Disease: Acc. Fed-IXI: Dice Fed-ISIC2019: Balanced Acc.
Global Local Global Local Global Local
FedAvg $0.724$ ($0.000$) $0.724$ ($0.000$) $0.9845^{*}$ ($0.0001$) $0.9846^{*}$ ($0.0001$) $0.657$ ($0.004$) $0.644$ ($0.004$)
FedAdam $0.719$ ($0.000$) $0.742$ ($0.000$) $0.9811^{*}$ ($0.0002$) $0.9811^{*}$ ($0.0002$) $0.647$ ($0.013$) $0.637$ ($0.012$)
FedProx $0.716$ ($0.000$) $0.721$ ($0.000$) $0.9842^{*}$ ($0.0001$) $0.9844^{*}$ ($0.0001$) $0.633$ ($0.022$) $0.652$ ($0.022$)
SCAFFOLD $0.711$ ($0.000$) $0.682$ ($0.000$) $0.9827^{*}$ ($0.0000$) $0.9827^{*}$ ($0.0000$) $\underline{0.672}^{*}$ ($0.013$) $0.669$ ($0.010$)
MOON $0.727$ ($0.010$) $0.752$ ($0.036$) $0.9852^{*}$ ($0.0002$) $0.9852^{*}$ ($0.0002$) $0.623$ ($0.021$) $0.605$ ($0.015$)
FedPer $\underline{0.814}^{*}$ ($0.002$) $0.9852^{*}$ ($0.0001$) $0.634$ ($0.014$)
Ditto $0.802^{*}$ ($0.005$) $0.9854^{*}$ ($0.0002$) $\mathbf{0.703}^{*}$ ($0.009$)
APFL $0.801^{*}$ ($0.006$) $\mathbf{0.9864}^{*}$ ($0.0002$) $0.608$ ($0.011$)
PerFCL $0.805^{*}$ ($0.018$) $0.9847^{*}$ ($0.0001$) $0.584$ ($0.019$)
FENDA-FL $\mathbf{0.815}^{*}$ ($0.000$) $\underline{0.9856}^{*}$ ($0.0001$) $0.607$ ($0.011$)
Silo $0.748$ ($0.000$) $0.9737$ ($0.0040$) $0.634$ ($0.025$)
Central $0.732$ ($0.000$) $0.9832$ ($0.0002$) $0.683$ ($0.020$)
Client 0 $0.735$ ($0.000$) $0.9648$ ($0.0123$) $0.549$ ($0.014$)
Client 1 $0.649$ ($0.000$) $0.9722$ ($0.0004$) $0.387$ ($0.029$)
Client 2 $0.589$ ($0.000$) $0.9550$ ($0.0015$) $0.554$ ($0.010$)
Client 3 $0.629$ ($0.000$) $0.444$ ($0.016$)
Client 4 $0.326$ ($0.018$)
Client 5 $0.285$ ($0.039$)
As seen in Table 2, each method produces a model that performs the Fed-IXI task well. However, there is a clear advantage to central training compared with individual client or siloed training. An even larger improvement is realized with FL. Among non-personalized FL methods, models trained with FedAvg, FedProx, and MOON surpass central training. All personalized FL approaches also improve upon the centrally trained model and produce four of the top five best metrics. APFL and FENDA-FL yield the best models, with APFL beating FENDA-FL by a small margin. Importantly, additional experiments, discussed in Appendix F, show that this gap can be closed with thoughtful architecture design for FENDA-FL. Finally, the performance variance for federally trained models is equivalent to the centrally trained version. As with Fed-Heart-Disease, using local checkpointing has small benefits for most methods.
Fed-ISIC2019 is a particularly challenging task. The results in Table 2 show that all FL strategies perform better than local models, with Ditto providing impressive results. It is the only method to surpass centralized training, though some other methods approach the performance. Most of the non-personalized FL methods equal or surpass siloed evaluation, though only SCAFFOLD’s model is statistically significant. Alternatively, for this task, all personalized FL approaches, with the stark exception of Ditto, under-perform. While the results are better than any of the local models, non-personalized methods, with the exception of MOON, consistently out-perform these techniques. Moreover, they fail to eclipse siloed accuracy. An investigation into where these approaches fall short for this dataset appears in Appendix G. It is worth noting that, for this task, Ditto required significantly more compute resources than any other method, but also achieves impressive accuracy.
FLamby Benchmark Improvements
Beyond the addition of personalized FL methods to the FLamby benchmark, the overall results achieved for Fed-IXI and FedISIC2019 stand in contrast to those reported in the original work. Therein, all FL strategies severely under-performed even local training for Fed-IXI. Similarly, none of the FL methods showed strong performance for Fed-ISIC2019 in the earlier study. This led the authors to conclude that federated training was not a useful approach for these tasks. The results presented here compellingly support the opposite conclusion.
In the original study, a standard SGD optimizer is used for client-side training, whereas AdamW is used here. Furthermore, the models for Fed-IXI and Fed-ISIC2019 employ Batch Normalization layers. When using FedAdam, it is critical that stats tracking be turned off for these layers to avoid issues during the aggregation process. This phenomenon is discussed in Appendix H. We speculate that a related issue may have impeded the convergence of the FedOpt family in their experiments.
5.1.2 GEMINI Results
For both GEMINI tasks, $50$ server rounds are performed. The clients in Mortality and Delirium prediction complete $2$ and $4$ local epochs, respectively, at each server round. GEMINI training setups, hyper-parameters, and other details are discussed in Appendix E. Models for each task are assessed based on accuracy and AUC/ROC.
@@ 5.1.2 GEMINI Results
Mean Accuracy Mean AUC/ROC
Global Local Global Local
FedAvg $0.8976$ ($0.0002$) $0.8971$ ($0.0006$) $0.8180^{*}$ ($0.0005$) $0.8177^{*}$ ($0.0005$)
FedAdam $0.8954$ ($0.0000$) $0.8954$ ($0.0000$) $0.8128$ ($0.0005$) $0.8123$ ($0.0008$)
FedProx $0.8957$ ($0.0004$) $0.8956$ ($0.0003$) $0.8174^{*}$ ($0.0004$) $0.8172^{*}$ ($0.0006$)
SCAFFOLD $0.8955$ ($0.0003$) $0.8954$ ($0.0000$) $\underline{0.8186}^{*}$ ($0.0002$) $0.8184^{*}$ ($0.0002$)
MOON $0.8968$ ($0.0008$) $0.8970$ ($0.0002$) $0.8184^{*}$ ($0.0003$) $0.8181^{*}$ ($0.0008$)
FedPer $\underline{0.8985}$ ($0.0005$) $0.8179^{*}$ ($0.0006$)
Ditto $\mathbf{0.8994}^{*}$ ($0.0003$) $\mathbf{0.8201}^{*}$ ($0.0011$)
APFL $0.8978$ ($0.0004$) $0.8138$ ($0.0017$)
PerFCL $0.8979$ ($0.0008$) $0.8117$ ($0.0012$)
FENDA-FL $0.8980$ ($0.0004$) $0.8181^{*}$ ($0.0001$)
Silo $0.8978$ ($0.0009$) $0.8140$ ($0.0012$)
Central $0.8981$ ($0.0005$) $0.8193$ ($0.0005$)
Client 0 $0.8966$ ($0.0014$) $0.8052$ ($0.0030$)
Client 1 $0.8930$ ($0.0025$) $0.7657$ ($0.0064$)
Client 2 $0.8943$ ($0.0060$) $0.7591$ ($0.0116$)
Client 3 $0.8954$ ($0.0000$) $0.7994$ ($0.0034$)
Client 4 $0.8955$ $(0.0005)$ $0.7974$ ($0.0037$)
Client 5 $0.8973$ $(0.0014)$ $0.7878$ ($0.0043$)
Client 6 $0.8954$ $(0.0000)$ $0.8019$ ($0.0035$)
@@
Mean Accuracy Mean AUC/ROC
Global Local Global Local
FedAvg $0.7987$ ($0.0127$) $0.8025$ ($0.0130$) $0.8302^{*}$ ($0.0146$) $0.8274^{*}$ ($0.0078$)
FedAdam $0.7689$ ($0.0036$) $0.7688$ ($0.0127$) $0.7897$ ($0.0078$) $0.7881$ ($0.0055$)
FedProx $\mathbf{0.8095}^{*}$ ($0.0036$) $0.8056$ ($0.0037$) $0.8488^{*}$ ($0.0043$) $\underline{0.8504}^{*}$ ($0.0042$)
SCAFFOLD $0.7480$ ($0.0000$) $0.7480$ ($0.0000$) $0.5491$ ($0.0469$) $0.5491$ ($0.0469$)
MOON $0.8056$ ($0.0100$) $0.8006$ ($0.0140$) $0.8375^{*}$ ($0.0051$) $0.8365^{*}$ ($0.0027$)
FedPer $\underline{0.8082}^{*}$ ($0.0024$) $0.8462^{*}$ ($0.0035$)
Ditto $0.7981$ ($0.0034$) $0.8271^{*}$ ($0.0017$)
APFL $0.8031$ ($0.0052$) $0.8430^{*}$ ($0.0062$)
PerFCL $0.7893$ ($0.0077$) $0.8111$ ($0.0089$)
FENDA-FL $0.8064$ ($0.0036$) $\mathbf{0.8518}^{*}$ ($0.0017$)
Silo $0.7936$ ($0.0102$) $0.8037$ ($0.0072$)
Central $0.8114$ ($0.0068$) $0.8458$ ($0.0026$)
Client 0 $0.7653$ ($0.0070$) $0.7977$ ($0.0128$)
Client 1 $0.7781$ ($0.0063$) $0.7795$ ($0.0161$)
Client 2 $0.7805$ ($0.0036$) $0.8185^{*}$ ($0.0021$)
Client 3 $0.7634$ ($0.0062$) $0.7539$ ($0.0055$)
Client 4 $0.7924$ ($0.0061$) $0.8179$ ($0.0076$)
Client 5 $0.8028$ ($0.0040$) $0.8429^{*}$ ($0.0018$)
@@
The results for the GEMINI Mortality task are shown in Table 3. FL methods improve upon locally trained models, reducing the performance gap to central training. This is especially true when considering AUC/ROC. In terms of accuracy, FedAvg is the only standard method that rivals the personalized FL algorithms and approaches siloed evaluation. On the other hand, most methods surpass the siloed AUC/ROC evaluation metric, nearing centrally trained model performance. Specifically, FedProx, SCAFFOLD, and MOON beat siloed training based on AUC/ROC, but did not in terms of accuracy. Ditto is the best approach, outperforming siloed and central accuracy and AUC/ROC. Considering both metrics, FedPer and FENDA-FL also provide strong results, each surpassing the siloed evaluation metrics. Both methods are competitive with non-personalized approaches in terms of AUC/ROC and outperform them based on accuracy.